You purchase 120 shares for $40 a share ($4,800), and after a year the price falls to $35. Calculate the percentage return on your investment if you bought the stock on margin and the margin requirement was (ignore commissions, dividends, and interest expense):

15 percent. Use a minus sign to enter the amount as a negative value. Round your answer to one decimal place.

55 percent. Use a minus sign to enter the amount as a negative value. Round your answer to one decimal place.

80 percent. Use a minus sign to enter the amount as a negative value. Round your answer to one decimal place.
Answer #1

Part 1:
Given that the margin required = 15%
Investment amount = $4800
Amount paid in the investment = (Investment amount)*(Margin required)= 4800*15% = $720
Sale amount = 120*35=$4200
Loss = Sale amount - Invested amount = 4200-4800 = -600
Loss (in %) = Loss/Amount paid in the investment = -600/720 = -83.33% or-83.3% (Rounded to one decimal place)

Part 2:
Given that the margin required = 55%
Investment amount = $4800
Amount paid in the investment = (Investment amount)*(Margin required)= 4800*55% = $2640
Sale amount = 120*35=$4200
Loss = Sale amount - Invested amount = 4200-4800 = -600
Loss (in %) = Loss/Amount paid in the investment = -600/2640 = -22.73% or -22.7% (Rounded to one decimal place)

Part 3:
Given that the margin required = 80%
Investment amount = $4800
Amount paid in the investment = (Investment amount)*(Margin required)= 4800*80% =$3840
Sale amount = 120*35=$4200
Loss = Sale amount - Invested amount = 4200-4800 = -600
Loss (in %) = Loss/Amount paid in the investment = -600/3840= -15.63% or -15.6% (Rounded to one decimal place)